# Flow-Aware Diffusion for Real-Time VR Restoration: Enhancing Spatiotemporal Coherence and Efficiency

### v1

@@ Introduction
Virtual Reality (VR) technologies have been widely adopted across various domains, including education, healthcare, industrial design, and entertainment [21, 18, 37]. However, one persistent challenge continues to hinder broader adoption: cybersickness. Commonly experienced as nausea, dizziness, or disorientation, this phenomenon remains a critical barrier to fully immersive experiences. Among its primary causes, the mismatch between visual and vestibular cues [26, 24], as described by the well-established sensory conflict theory [39], has been widely recognized. This mismatch is particularly pronounced in visually rich environments, where artificial motion cues dominate the user’s field of view. In such scenarios, visual motion alone can evoke a strong illusion of self-motion, even in the absence of physical movement.
A major source of this sensory mismatch is unnatural or excessive visual motion, which can induce discomfort even when users are physically stationary. In particular, optical flow, which refers to the perceived motion of the visual scene relative to the observer [30], has been identified as a key visual trigger. High-intensity optical flow creates a compelling illusion of self-motion; when unaccompanied by corresponding vestibular input, it amplifies perceptual inconsistencies and leads to cybersickness. This effect is especially pronounced in peripheral regions of the visual field, where motion sensitivity is heightened. As such, modulating the structure and magnitude of optical flow in rendered content [45, 28] has become a critical strategy for enhancing comfort across a wide range of immersive applications—not limited to virtual locomotion, but including diverse interactive and head-tracked experiences.
A variety of hardware and software strategies have been proposed to mitigate cybersickness, including vestibular stimulation and visual content adaptation. Hardware-based methods such as Galvanic Vestibular Stimulation (GVS)[42] and motion platforms[29] aim to align physical and visual cues but are often intrusive and impractical for consumer use. Software approaches, including virtual visual anchors[47], dynamic field-of-view restriction[10], and peripheral blurring [6], modify rendered content to reduce perceived motion, though they can compromise immersion or visual quality. Notably, Lou et al. introduced a geometric simplification framework that reduces peripheral optical flow by manually editing high-motion regions of the virtual scene [33, 34]. Their method segments scene geometry based on offline optical flow analysis and replaces complex elements with simplified shapes. While effective, it relies heavily on predefined meshes, manual tuning, and offline and asynchronous preprocessing, limiting its scalability in dynamic or procedurally generated environments.
@@ 1.1 Cybersickness and Visual Motion: A Sensory Conflict Perspective
As VR technologies become increasingly prevalent in education, healthcare, and entertainment, one of the major barriers to mainstream adoption remains cybersickness, which is a form of motion-induced discomfort typically experienced as nausea, dizziness, or disorientation [25, 43]. A widely accepted explanation for cybersickness is the sensory conflict theory, which attributes user discomfort to mismatches between visual and vestibular inputs [39]. When users perceive motion visually, for instance through large field-of-view displays or scene movement, a lack of corresponding physical feedback can result in a conflict. This conflict often leads to perceptual instability.
In immersive VR environments, this conflict is often intensified by dynamic visual stimuli that produce strong optical flow, particularly in peripheral vision, where motion sensitivity is heightened [30]. Excessive or inconsistent optical flow can induce a compelling illusion of self-motion, aggravating sensory mismatch, and triggering discomfort. Moreover, visual artifacts such as frame skipping, ghosting, and temporal discontinuity further disrupt motion continuity and exacerbate user disorientation [43].
Given the real-time and high-resolution demands of modern VR systems often require high frame rates [22, 19], so that cybersickness remains a critical challenge. Addressing it requires not only improving video quality but also ensuring perceptual coherence between the visual stimulus and the user’s motion expectations. As such, motion-consistent rendering and optical flow suppression have become essential goals in optimizing VR content.
@@ 1.2 Cybersickness and optical flow
Temporal consistency is essential for comfortable video perception, particularly in immersive VR environments where users are highly sensitive to motion discontinuities. One of the most critical visual cues contributing to motion perception and even can be extended to cybersickness is optical flow, which encodes pixel-wise motion between consecutive frames [30]. Excessive or inconsistent optical flow, especially in peripheral vision, can induce the illusion of self-motion and lead to a sensory conflict when not accompanied by corresponding vestibular input [43]. This mismatch is a well-established cause of cybersickness [39], and managing optical flow has thus become a central concern in perceptual stability research for VR.
From a modeling perspective, optical flow has been widely used as a prior to improve temporal continuity in video restoration tasks [36, 8]. These methods perform well in scenarios involving fine-grained motion; however, their reliance on pre-computed or external flow estimates limits their ability to adapt to varied and rapidly changing motion patterns dynamically. Moreover, traditional flow-based models typically treat motion information as an auxiliary signal [16, 49] rather than a central learning objective. This often leads to suboptimal integration between motion cues and content reconstruction. Their architectures are rarely end-to-end trainable, which further restricts flexibility in high-frame-rate, wide-field VR applications where temporal coherence is paramount.
Recent advances in generative modeling, particularly diffusion-based frameworks, have begun to explore the integration of motion cues into the generation process. For example, Control-A-Video [7] introduces flow-conditioned denoising to improve motion controllability and visual stability in text-to-video synthesis. However, most existing models still lack explicit motion supervision, making them ill-suited for scenarios where motion fidelity is crucial, such as VR where even subtle inconsistencies between expected and observed motion can disrupt vestibular-visual alignment and induce disorientation.
@@ 1.3 Explicit Geometry-Based Flow Reduction
One direct strategy for reducing cybersickness is to explicitly suppress optical flow through geometric manipulation of the visual scene. Such methods modify the structure or appearance of virtual content to reduce perceived motion, particularly in peripheral regions of the field of view where motion sensitivity is greatest [30]. In general, these approaches can be divided into two categories: scene simplification and visual anchoring.
Scene simplification methods attempt to lower motion saliency by reducing visual complexity in high-flow areas. Lou et al. [33] proposed a framework that segments a scene based on optical flow magnitude and applies geometric simplification to the high-motion regions, typically in the periphery. Their follow-up work [34] rendered original and simplified scenes in concentric fields of view, preserving fidelity while suppressing disruptive motion. Similar efforts include deforming peripheral scene geometry to align with user motion direction, thereby minimizing perceived motion [5, 44]. Visual anchoring techniques, on the contrary, aim to reduce the sensation of motion by introducing static visual references. A classic example is the insertion of a virtual nose into the user’s field of view, which serves as a stable reference point and helps ground the user’s perception [47]. Another strategy is peripheral blurring during rapid motion events, which limits visual flow without interfering with central vision [6].
@@ 1.4 Learning-Based Flow-Aware Modeling
In contrast to explicit manipulation of scene geometry, learning-based methods seek to model and regulate motion perception through data-driven optimization. These approaches aim to internalize motion continuity and generate temporally coherent content that minimizes perceptual conflicts without manual intervention. Early flow-aware learning models, such as DAIN [3] and FLAVR [20], incorporate pre-computed optical flow to warp adjacent frames, reducing artifacts like ghosting and flicker. However, these methods treat flow as a secondary alignment signal, limiting adaptability and robustness in immersive settings. Moreover, they lack end-to-end differentiability, which restricts their generalization to unpredictable motion patterns or scene dynamics common in VR.
Recent advancements explore deeper integration of motion priors into generative architectures. Transformer-based models like VRT [31] have demonstrated the ability to capture long-range temporal dependencies, enhancing consistency across time. However, their high computational demands and lack of explicit perceptual supervision make them unsuitable for real-time, high-frame-rate applications such as VR. Diffusion-based generative models have recently emerged as a promising alternative, offering both quality and flexibility. Control-A-Video [7] conditions denoising flow features to produce controllable, temporally stable sequences. However, even these models often omit explicit motion supervision, making them vulnerable to subtle but perceptionally significant inconsistencies.
@@ 2 Methodology
To more efficiently modulate the structure and magnitude of optical flow so as to reduce the cybersickness in real-time , we propose the U-shaped Mamba Diffusion (U-MAD) framework, which aims to restore high-quality video sequences from spatially and temporally degraded inputs. The framework, shown in Fig. 1, takes as input a sequence of degraded frames $F_{\text{deg}}=\{f_{1},f_{2},\dots,f_{T}\}$ and a downsampled clean reference $F_{\text{raw}}$, available during training. The degraded frames are first spatially cropped into patches to improve processing efficiency, then passed through a U-shaped encoder–decoder backbone built on Mamba [12], a state-space sequence model that enables scalable modeling of long-range temporal dependencies.
@@ 2.2 U-shaped Mamba Diffusion Backbone
The core of our architecture is a U-shaped diffusion network composed of hierarchical Mamba-based encoder and decoder blocks. Instead of traditional CNN or Transformer layers, we adopt Mamba [12], a state-space sequence model that excels in modeling long-range temporal dependencies with linear computational complexity. Our implementation follows the recent advancements from Structured State Space Sequence Models (S4) [13] to the more selective and efficient u-Mamba (S6) [38].
Each Mamba block operates over spatio-temporal feature maps with shape $(B,C,H,W,D)$, where $B$ is the batch size, $C$ is the number of channels, $(H,W)$ are spatial dimensions, and $D$ is the temporal depth. The input features are normalized and transposed as necessary before being passed into the Mamba structure. Internally, each block includes two parallel branches designed to model complementary aspects of the input: The first branch captures global sequence features through pure linear transformation and activation, and the The second branch enhances local spatial awareness by applying convolution and a global-local attention mechanism (GLAM) [32]:
[…]
@@
Finally, the denoising diffusion process is embedded within this U-shaped backbone. At each timestep $t$, a noisy representation $x_{t}$ is iteratively refined by the network toward the clean latent code $x_{0}$, under a temporally structured DDPM formulation [48]. The detailed mechanism of conditional flow guidance is discussed in the following section.
@@ 2.3 Unified Motion-Structure Embedding (UMSE)
[…]
@@
In addition to motion cues, UMSE incorporates structural priors inspired by LOST [9], including frame index, window scale, and encoding quality parameters. These factors are first normalized and embedded into a compact vector using dedicated MLPs, then spatially broadcasted to match the resolution of the flow features.
@@ 2.4 Training Objective
  • Reconstruction Loss $\mathcal{L}_{\text{rec}}$: We use the Charbonnier loss [4] to robustly measure pixel-level differences between output and ground truth:
@@ 3.1.1 Datasets
The first dataset we use for training consists of 19 stereoscopic VR videos, as provided in [40]. The video sizes and lengths vary from one another. The width $\times$ height size of 9 videos is 2880 $\times$ 1080, while that of 8 videos is 1920 $\times$ 720, and the remaining two videos are 2730 $\times$ 1024. There are four videos of 61 seconds, four videos of 60.5 seconds, and 11 videos of 60 seconds. To test the robustness of our proposed model, we take another dataset of SEPE8K videos [1]. Both datasets’ information is given in Table 1.
@@ 3.3.1 Backbone Architecture Comparison
  • U-Net Diffusion: a 3D convolutional U-Net as proposed in VideoDiffusion [17]
  • Transformer Diffusion: a DiT-style diffusion model adapted to video frames [41]
@@ 3.3.2 Conditioning Strategy Comparison
Many recent studies have leveraged camera movement or other signals to guide video restoration. For instance, MotionCtrl [46] and CameraCtrl [15] utilize camera motion as a control signal. However, one of the key factors contributing to cybersickness is the inconsistency in motion between frames. To address this, we incorporated optical flow consistency into our model, which helps maintain smoother motion and reduces the likelihood of inducing user discomfort.
@@
Motion Vectors [46] 30.9 0.892 0.135 0.115
Camera Params [15] 30.1 0.886 0.142 0.123
@@ 3.4 Domain-specific Effectiveness and Generalization
To evaluate how well our method generalizes in different video domains, we conduct experiments on both conventional forward-view datasets and VR-specific omnidirectional datasets. These domains differ significantly in motion characteristics, field-of-view (FOV), and spatial continuity. Conventional 2D video datasets (e.g., SEPE8k [1]) contain standard perspective videos, whereas VR datasets, like Padmanaban [40], feature wide-field scenes with strong camera and object motion. This setup allows us to examine both generalization and domain-specific effectiveness. Table 4 summarizes the performance of our model in both domains. While our method performs well in conventional videos, its advantage becomes more pronounced in VR settings, particularly in LPIPS and tOF. These results indicate that motion-aware conditioning and Mamba-based temporal modeling are especially beneficial in wide-field, high-motion environments typical of VR content.

### v2

@@ Introduction
Virtual Reality (VR) technologies have been widely adopted across various domains, including education, healthcare, industrial design, and entertainment [kaminska2019virtual, 10.1145/3680533.3697068, luo2024using]. However, one persistent challenge continues to hinder broader adoption: cybersickness. Commonly experienced as nausea, dizziness, or disorientation, this phenomenon remains a critical barrier to fully immersive experiences. Among its primary causes, the mismatch between visual and vestibular cues [Kim2020ADeepMotionSickness, keshavarz2015vection], as described by the well-established sensory conflict theory [Nürnberger2021MismatchofVisual-Vestibular], has been widely recognized. This mismatch is particularly pronounced in visually rich environments, where artificial motion cues dominate the user’s field of view. In such scenarios, visual motion alone can evoke a strong illusion of self-motion, even in the absence of physical movement.
A major source of this sensory mismatch is unnatural or excessive visual motion, which can induce discomfort even when users are physically stationary. In particular, optical flow, which refers to the perceived motion of the visual scene relative to the observer [lefevre2009optical], has been identified as a key visual trigger. High-intensity optical flow creates a compelling illusion of self-motion; when unaccompanied by corresponding vestibular input, it amplifies perceptual inconsistencies and leads to cybersickness. This effect is especially pronounced in peripheral regions of the visual field, where motion sensitivity is heightened. As such, modulating the structure and magnitude of optical flow in rendered content [sun2024pixel, kreinin2023highfidelity] has become a critical strategy for enhancing comfort across a wide range of immersive applications—not limited to virtual locomotion, but including diverse interactive and head-tracked experiences.
A variety of hardware and software strategies have been proposed to mitigate cybersickness, including vestibular stimulation and visual content adaptation. Hardware-based methods such as Galvanic Vestibular Stimulation (GVS)[pradhan2022visual] and motion platforms[langbehn2017application] aim to align physical and visual cues but are often intrusive and impractical for consumer use. Software approaches, including virtual visual anchors[whittinghill2015nasum], dynamic field-of-view restriction[fernandes2016combating], and peripheral blurring [budhiraja2017rotation], modify rendered content to reduce perceived motion, though they can compromise immersion or visual quality. Notably, Lou et al. introduced a geometric simplification framework that reduces peripheral optical flow by manually editing high-motion regions of the virtual scene [lou2022general, lou2022geometric]. Their method segments scene geometry based on offline optical flow analysis and replaces complex elements with simplified shapes. While effective, it relies heavily on predefined meshes, manual tuning, and offline and asynchronous preprocessing, limiting its scalability in dynamic or procedurally generated environments.
@@ 1.1 Cybersickness and Visual Motion: A Sensory Conflict Perspective
As VR technologies become increasingly prevalent in education, healthcare, and entertainment, one of the major barriers to mainstream adoption remains cybersickness, which is a form of motion-induced discomfort typically experienced as nausea, dizziness, or disorientation [Kim2020Multisensory, Regan2017The]. A widely accepted explanation for cybersickness is the sensory conflict theory, which attributes user discomfort to mismatches between visual and vestibular inputs [Nürnberger2021MismatchofVisual-Vestibular]. When users perceive motion visually, for instance through large field-of-view displays or scene movement, a lack of corresponding physical feedback can result in a conflict. This conflict often leads to perceptual instability.
In immersive VR environments, this conflict is often intensified by dynamic visual stimuli that produce strong optical flow, particularly in peripheral vision, where motion sensitivity is heightened [lefevre2009optical]. Excessive or inconsistent optical flow can induce a compelling illusion of self-motion, aggravating sensory mismatch, and triggering discomfort. Moreover, visual artifacts such as frame skipping, ghosting, and temporal discontinuity further disrupt motion continuity and exacerbate user disorientation [Regan2017The].
Given the real-time and high-resolution demands of modern VR systems often require high frame rates [kemeny2020getting, jung2021floor], so that cybersickness remains a critical challenge. Addressing it requires not only improving video quality but also ensuring perceptual coherence between the visual stimulus and the user’s motion expectations. As such, motion-consistent rendering and optical flow suppression have become essential goals in optimizing VR content.
@@ 1.2 Cybersickness and optical flow
Temporal consistency is essential for comfortable video perception, particularly in immersive VR environments where users are highly sensitive to motion discontinuities. One of the most critical visual cues contributing to motion perception and even can be extended to cybersickness is optical flow, which encodes pixel-wise motion between consecutive frames [lefevre2009optical]. Excessive or inconsistent optical flow, especially in peripheral vision, can induce the illusion of self-motion and lead to a sensory conflict when not accompanied by corresponding vestibular input [Regan2017The]. This mismatch is a well-established cause of cybersickness [Nürnberger2021MismatchofVisual-Vestibular], and managing optical flow has thus become a central concern in perceptual stability research for VR.
From a modeling perspective, optical flow has been widely used as a prior to improve temporal continuity in video restoration tasks [lu2022priors, cheng2017segflow]. These methods perform well in scenarios involving fine-grained motion; however, their reliance on pre-computed or external flow estimates limits their ability to adapt to varied and rapidly changing motion patterns dynamically. Moreover, traditional flow-based models typically treat motion information as an auxiliary signal [henter2020moglow, zand2023flow] rather than a central learning objective. This often leads to suboptimal integration between motion cues and content reconstruction. Their architectures are rarely end-to-end trainable, which further restricts flexibility in high-frame-rate, wide-field VR applications where temporal coherence is paramount.
Recent advances in generative modeling, particularly diffusion-based frameworks, have begun to explore the integration of motion cues into the generation process. For example, Control-A-Video [chen2024controlavideocontrollabletexttovideodiffusion] introduces flow-conditioned denoising to improve motion controllability and visual stability in text-to-video synthesis. However, most existing models still lack explicit motion supervision, making them ill-suited for scenarios where motion fidelity is crucial, such as VR where even subtle inconsistencies between expected and observed motion can disrupt vestibular-visual alignment and induce disorientation.
@@ 1.3 Explicit Geometry-Based Flow Reduction
One direct strategy for reducing cybersickness is to explicitly suppress optical flow through geometric manipulation of the visual scene. Such methods modify the structure or appearance of virtual content to reduce perceived motion, particularly in peripheral regions of the field of view where motion sensitivity is greatest [lefevre2009optical]. In general, these approaches can be divided into two categories: scene simplification and visual anchoring.
Scene simplification methods attempt to lower motion saliency by reducing visual complexity in high-flow areas. Lou et al. [lou2022general] proposed a framework that segments a scene based on optical flow magnitude and applies geometric simplification to the high-motion regions, typically in the periphery. Their follow-up work [lou2022geometric] rendered original and simplified scenes in concentric fields of view, preserving fidelity while suppressing disruptive motion. Similar efforts include deforming peripheral scene geometry to align with user motion direction, thereby minimizing perceived motion [bruder2011tuning, so2001metric]. Visual anchoring techniques, on the contrary, aim to reduce the sensation of motion by introducing static visual references. A classic example is the insertion of a virtual nose into the user’s field of view, which serves as a stable reference point and helps ground the user’s perception [whittinghill2015nasum]. Another strategy is peripheral blurring during rapid motion events, which limits visual flow without interfering with central vision [budhiraja2017rotation].
@@ 1.4 Learning-Based Flow-Aware Modeling
In contrast to explicit manipulation of scene geometry, learning-based methods seek to model and regulate motion perception through data-driven optimization. These approaches aim to internalize motion continuity and generate temporally coherent content that minimizes perceptual conflicts without manual intervention. Early flow-aware learning models, such as DAIN [bao2019depth] and FLAVR [kalluri2023flavr], incorporate pre-computed optical flow to warp adjacent frames, reducing artifacts like ghosting and flicker. However, these methods treat flow as a secondary alignment signal, limiting adaptability and robustness in immersive settings. Moreover, they lack end-to-end differentiability, which restricts their generalization to unpredictable motion patterns or scene dynamics common in VR.
Recent advancements explore deeper integration of motion priors into generative architectures. Transformer-based models like VRT [liang2022recurrent] have demonstrated the ability to capture long-range temporal dependencies, enhancing consistency across time. However, their high computational demands and lack of explicit perceptual supervision make them unsuitable for real-time, high-frame-rate applications such as VR. Diffusion-based generative models have recently emerged as a promising alternative, offering both quality and flexibility. Control-A-Video [chen2024controlavideocontrollabletexttovideodiffusion] conditions denoising flow features to produce controllable, temporally stable sequences. However, even these models often omit explicit motion supervision, making them vulnerable to subtle but perceptionally significant inconsistencies.
@@ 2 Methodology
To more efficiently modulate the structure and magnitude of optical flow so as to reduce the cybersickness in real-time , we propose the U-shaped Mamba Diffusion (U-MAD) framework, which aims to restore high-quality video sequences from spatially and temporally degraded inputs. The framework, shown in Fig. 1, takes as input a sequence of degraded frames $F_{\text{deg}}=\{f_{1},f_{2},\dots,f_{T}\}$ and a downsampled clean reference $F_{\text{raw}}$, available during training. The degraded frames are first spatially cropped into patches to improve processing efficiency, then passed through a U-shaped encoder–decoder backbone built on Mamba [gu2024mambalineartimesequencemodeling], a state-space sequence model that enables scalable modeling of long-range temporal dependencies.
@@ 2.2 U-shaped Mamba Diffusion Backbone
The core of our architecture is a U-shaped diffusion network composed of hierarchical Mamba-based encoder and decoder blocks. Instead of traditional CNN or Transformer layers, we adopt Mamba [gu2024mambalineartimesequencemodeling], a state-space sequence model that excels in modeling long-range temporal dependencies with linear computational complexity. Our implementation follows the recent advancements from Structured State Space Sequence Models (S4) [gu2022efficientlymodelinglongsequences] to the more selective and efficient u-Mamba (S6) [ma2024umambaenhancinglongrangedependency].
Each Mamba block operates over spatio-temporal feature maps with shape $(B,C,H,W,D)$, where $B$ is the batch size, $C$ is the number of channels, $(H,W)$ are spatial dimensions, and $D$ is the temporal depth. The input features are normalized and transposed as necessary before being passed into the Mamba structure. Internally, each block includes two parallel branches designed to model complementary aspects of the input: The first branch captures global sequence features through pure linear transformation and activation, and the The second branch enhances local spatial awareness by applying convolution and a global-local attention mechanism (GLAM) [liu2021globalattentionmechanismretain]:
$\text{Branch 1: }X_{1}=\text{SiLU}(\text{Linear}(\text{LayerNorm}(X)))$
$\text{Branch 2: }X_{2}=\text{GLAM}(\text{SiLU}(\text{Conv}(\text{Linear}(\text{LayerNorm}(X)))))$
@@
Finally, the denoising diffusion process is embedded within this U-shaped backbone. At each timestep $t$, a noisy representation $x_{t}$ is iteratively refined by the network toward the clean latent code $x_{0}$, under a temporally structured DDPM formulation [yang2024survey]. The detailed mechanism of conditional flow guidance is discussed in the following section.
@@ 2.3 Unified Motion-Structure Embedding (UMSE)
where $S\in\mathbb{R}^{H\times W\times H\times W}$ and $\langle\cdot,\cdot\rangle$ denotes the dot product. These correlations are used to iteratively refine an optical flow field $\mathbf{u}_{t}$ through a learned residual predictor:
@@
In addition to motion cues, UMSE incorporates structural priors inspired by LOST [dehaghi2024reversingdamageqpawaretransformerdiffusion], including frame index, window scale, and encoding quality parameters. These factors are first normalized and embedded into a compact vector using dedicated MLPs, then spatially broadcasted to match the resolution of the flow features.
@@ 2.4 Training Objective
  • Reconstruction Loss $\mathcal{L}_{\text{rec}}$: We use the Charbonnier loss [barron2019general] to robustly measure pixel-level differences between output and ground truth:
@@ 3.1.1 Datasets
The first dataset we use for training consists of 19 stereoscopic VR videos, as provided in [padmanaban2018TowardsaMachineLearningApproach]. The video sizes and lengths vary from one another. The width $\times$ height size of 9 videos is 2880 $\times$ 1080, while that of 8 videos is 1920 $\times$ 720, and the remaining two videos are 2730 $\times$ 1024. There are four videos of 61 seconds, four videos of 60.5 seconds, and 11 videos of 60 seconds. To test the robustness of our proposed model, we take another dataset of SEPE8K videos [Al2023SEPEDataset]. Both datasets’ information is given in Table 1.
@@ 3.3.1 Backbone Architecture Comparison
  • U-Net Diffusion: a 3D convolutional U-Net as proposed in VideoDiffusion [ho2022videodiffusionmodels]
  • Transformer Diffusion: a DiT-style diffusion model adapted to video frames [peebles2023scalablediffusionmodelstransformers]
@@ 3.3.2 Conditioning Strategy Comparison
Many recent studies have leveraged camera movement or other signals to guide video restoration. For instance, MotionCtrl [wang2024motionctrlunifiedflexiblemotion] and CameraCtrl [he2024cameractrlenablingcameracontrol] utilize camera motion as a control signal. However, one of the key factors contributing to cybersickness is the inconsistency in motion between frames. To address this, we incorporated optical flow consistency into our model, which helps maintain smoother motion and reduces the likelihood of inducing user discomfort.
@@
Motion Vectors [wang2024motionctrlunifiedflexiblemotion] 30.9 0.892 0.135 0.115
Camera Params [he2024cameractrlenablingcameracontrol] 30.1 0.886 0.142 0.123
@@ 3.4 Domain-specific Effectiveness and Generalization
To evaluate how well our method generalizes in different video domains, we conduct experiments on both conventional forward-view datasets and VR-specific omnidirectional datasets. These domains differ significantly in motion characteristics, field-of-view (FOV), and spatial continuity. Conventional 2D video datasets (e.g., SEPE8k [Al2023SEPEDataset]) contain standard perspective videos, whereas VR datasets, like Padmanaban [padmanaban2018TowardsaMachineLearningApproach], feature wide-field scenes with strong camera and object motion. This setup allows us to examine both generalization and domain-specific effectiveness. Table 4 summarizes the performance of our model in both domains. While our method performs well in conventional videos, its advantage becomes more pronounced in VR settings, particularly in LPIPS and tOF. These results indicate that motion-aware conditioning and Mamba-based temporal modeling are especially beneficial in wide-field, high-motion environments typical of VR content.